OpenStudy (anonymous):

How does one get from sin(π/4) to (√2/2)? As in what's the rule?

OpenStudy (lgbasallote):

are you familiar that \(\pi / 4\) = 45 degrees?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

this is a picture of a 45-45-90 triangle  if you look at that..sin (45) = opposite/hypotenuse = \(1/\sqrt 2\)

OpenStudy (lgbasallote):

if i rationalize that i get \[\frac{1}{\sqrt 2} \times \frac{\sqrt 2}{\sqrt 2} = \frac{\sqrt 2}{2}\] got it?

OpenStudy (anonymous):

so how does sin405 ° end up being (√2/2) trigonometry

OpenStudy (anonymous):

without a calculator. as in using the fact that the trigonometric functions are periodic

OpenStudy (callisto):

sin405 = sin (360 + 45) = sin 45

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

welcome~